# Theory of cell size regulation during migration in adhered cells

### v1

@@ Abstract
model- We consider a simplified model of a cell adhering to an underlying substrate and adopting the shape of a spherical cap (see Fig.1). The cell interior is modeled as a solution, comprising of water, ions that exchange with the extracellular environment, and impermeant molecules. Among the impermeant molecules are proteins that interact with cortical actin and act as polarity cues (see Fig. 1).
@@
where $p_{n}$ are polynomial coefficients and where even powers have vanished due to the angular integration. The cell size enters Eq. (Theory of cell size regulation during migration in adhered cells) through the radius $R$ and consequently affects the migration velocity. Its effect is two-fold: (i) It enters the prefactor on the left-hand side and acts in similar to an inverse temperature near a phase transition, and (ii) it defines the typical magnitude of migration speed according to $u=\tilde{u}D/R$.
@@
For $p_{3}<0$, we neglect $p_{5}$ and Eq. (Theory of cell size regulation during migration in adhered cells) reduces to a quadratic equation for $\tilde{u}$. We solve it (see Appendix C) and find that $\tilde{u}=0$ for small radii and smoothly increases $\tilde{u}\geq 0$ beyond the critical radius $R_{c}=\Gamma D/f_{0}p_{1}.$ This smooth transition corresponds to a supercritical bifurcation when velocity dynamics are accounted (Fig. 3 and Appendix C). For $p_{3}>0$, Eq. (Theory of cell size regulation during migration in adhered cells) reduces to a quadratic equation for $\tilde{u}^{2}$. Its solution is $\tilde{u}=0$ for $R<R_{c}^{\ast},$ with $R_{c}^{*}=R_{c}/\left(1+p_{3}^{2}/4p_{1}|p_{5}|\right)<R_{c},$ above which there is a sudden jump to $\tilde{u}=u_{c}^{\ast}R/D>0.$ This discontinuous transition corresponds to a subcritical bifurcation (Fig. 3 and Appendix C). In both cases, $\tilde{u}$ increases with $R.$ However, as $u=\tilde{u}D/R,$ the actual velocity exhibits a maximum and decreases with $R$ for large radii, as is shown in Fig. 3. Explicitly, the velocity is given by
@@
Comparison with experiments - We compare our theoretical predictions with recent experiments by Nagy et al. [22] on neutrophil swelling and migration. In each experiment, several hundred adhered neutrophils were simultaneously and uniformly exposed to chemoattractants. For each cell, velocity was inferred from time-lapse tracking, contact area was measured directly, and volume was quantified using Fluorescence Exclusion Microscopy [22, 41]. We begin by summarizing the main experimental findings and interpret them through the lens of our theory. As shown in Fig. 4, our predictions show excellent agreement with the measured data.
Cell spreading and swelling. The cells exhibit two phases in their response to chemoattractant exposure: spreading and swelling (see Fig. 4). The cell spreads within a couple of minutes, where the contact area increases by about $10\%$, while the volume decreases by around $5\%$. In the following $15$ minutes, the cell swells. The area continues to increase by an additional $20\%$, while the volume increases to approximately $15\%$ more than the initial volume. The swelling was shown to be associated with ion transport and was suppressed upon inhibition of NHE1 exchangers [22], which play an important role in regulatory volume increase  [28, 20, 21].
@@
Cell migration. The experiments also reveal an interesting behavior of the cellular velocities. Before the exposure to chemo-attractants, the cells move randomly with a typical velocity of about 2 µm/min [22]. This motion corresponds to a random velocity $u_{\xi}$ that is possible even for $u=0$, due to fluctuations in the active force, as discussed above. Upon exposure, the velocity increases and reaches a new typical value, about twice as large as the original one (see Fig. 4c), which remains relatively constant on the scale of $10$ minutes. During this time, the cellular motion remains random with low persistence. We regard this as a change in $u_{\xi}$, whose modeling lies beyond the scope of this work.
@@
Discussion - This Letter introduces a minimal framework to elucidate cellular size regulation during migration, predicting distinct regimes of migration velocity and its dependence on cell size, including a critical contact area for migration. Our predictions can be tested experimentally by simultaneously tracking cell position, contact area, and volume across various cell types.
@@
Acknowledgments. We thank Tamas L. Nagy, Nir S. Gov, Samuel A. Safran and Kinneret Keren for useful discussions and suggestions.

### v2

@@ I Introduction
II The model
We consider a simplified model of a cell adhering to an underlying substrate and adopting the shape of a spherical cap (see Fig.1). The cell interior is modeled as a solution, comprising of water, ions that exchange with the extracellular environment, and impermeant molecules. Among the impermeant molecules are proteins that interact with cortical actin and act as polarity cues (see Fig. 1).
@@ II The model
where $p_{n}$ are polynomial coefficients and where even powers have vanished due to the angular integration. The cell size enters Eq. (II) through the radius $R$ and consequently affects the migration velocity. Its effect is two-fold: (i) It enters the prefactor on the left-hand side and acts in similar to an inverse temperature near a phase transition, and (ii) it defines the typical magnitude of migration speed according to $u=\tilde{u}D/R$.
@@
For $p_{3}<0$, we neglect $p_{5}$ and Eq. (II) reduces to a quadratic equation for $\tilde{u}$. We solve it (see Appendix C) and find that $\tilde{u}=0$ for small radii and smoothly increases $\tilde{u}\geq 0$ beyond the critical radius $R_{c}=\Gamma D/f_{0}p_{1}.$ This smooth transition corresponds to a supercritical bifurcation when velocity dynamics are accounted (Fig. 3 and Appendix C). For $p_{3}>0$, Eq. (II) reduces to a quadratic equation for $\tilde{u}^{2}$. Its solution is $\tilde{u}=0$ for $R<R_{c}^{\ast},$ with $R_{c}^{*}=R_{c}/\left(1+p_{3}^{2}/4p_{1}|p_{5}|\right)<R_{c},$ above which there is a sudden jump to $\tilde{u}=u_{c}^{\ast}R/D>0.$ This discontinuous transition corresponds to a subcritical bifurcation (Fig. 3 and Appendix C). In both cases, $\tilde{u}$ increases with $R.$ However, as $u=\tilde{u}D/R,$ the actual velocity exhibits a maximum and decreases with $R$ for large radii, as is shown in Fig. 3. Explicitly, the velocity is given by
@@
III Comparison with experiments
We compare our theoretical predictions with recent experiments by Nagy et al. [22] on neutrophil swelling and migration. In each experiment, several hundred adhered neutrophils were simultaneously and uniformly exposed to chemoattractants. For each cell, velocity was inferred from time-lapse tracking, contact area was measured directly, and volume was quantified using Fluorescence Exclusion Microscopy [22, 41]. We begin by summarizing the main experimental findings and interpret them through the lens of our theory. As shown in Fig. 4, our predictions show excellent agreement with the measured data.
Cell spreading and swelling - The cells exhibit two phases in their response to chemoattractant exposure: spreading and swelling (see Fig. 4). The cell spreads within a couple of minutes, where the contact area increases by about $10\%$, while the volume decreases by around $5\%$. In the following $15$ minutes, the cell swells. The area continues to increase by an additional $20\%$, while the volume increases to approximately $15\%$ more than the initial volume. The swelling was shown to be associated with ion transport and was suppressed upon inhibition of NHE1 exchangers [22], which play an important role in regulatory volume increase  [28, 20, 21].
@@ III Comparison with experiments
Cell migration - The experiments also reveal an interesting behavior of the cellular velocities. Before the exposure to chemo-attractants, the cells move randomly with a typical velocity of about 2 µm/min [22]. This motion corresponds to a random velocity $u_{\xi}$ that is possible even for $u=0$, due to fluctuations in the active force, as discussed above. Upon exposure, the velocity increases and reaches a new typical value, about twice as large as the original one (see Fig. 4c), which remains relatively constant on the scale of $10$ minutes. During this time, the cellular motion remains random with low persistence. We regard this as a change in $u_{\xi}$, whose modeling lies beyond the scope of this work.
@@
IV Discussion
This Letter introduces a minimal framework to elucidate cellular size regulation during migration, predicting distinct regimes of migration velocity and its dependence on cell size, including a critical contact area for migration. Our predictions can be tested experimentally by simultaneously tracking cell position, contact area, and volume across various cell types.
@@ IV Discussion
V Acknowledgments.
We thank Tamas L. Nagy, Nir S. Gov, Samuel A. Safran and Kinneret Keren for useful discussions and suggestions.